OpenStudy (anonymous):

What is the slope of the tangent line passing through the point (2, 3) of the graph of this function? f(x) = x2 + 2x − 5 1 2 3 4 6

OpenStudy (michele_laino):

hint: we have to compute the first derivative of f(x) at point x=2

OpenStudy (anonymous):

f(x) =3

OpenStudy (michele_laino):

no, we have to compute the first derivative of f(x): \[f'\left( x \right) = ...?\]

OpenStudy (michele_laino):

hint: the first derivative of x^ is 2*x and the first derivative of x, is 1

OpenStudy (michele_laino):

oops..x^2

OpenStudy (anonymous):

x^ is x^2?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

so the first derivative of f(x), is equal to the first derivative of x^2 plus the first derivative of 2x please note that the first derivative of -5 is zero

OpenStudy (anonymous):

so its 1

OpenStudy (michele_laino):

the first derivative of f(x), is: f ' (x)=2x+2 so, what is f ' (2)=...?

OpenStudy (michele_laino):

f ' (2) = 2*2+2=...?

OpenStudy (anonymous):

6

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ok I had got 6 but I was second guessing got confused

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!